The risk free rate of return is 1.8% and the expected return on the market portfolio is 8.35%. Given the following possible returns for Lidar Limited and the S&P/TSX Composite Index(found in Table 1.1):

a.Calculate the expected return for Lidar (illustrate your solution using MS Equation Editor).

b.Calculate the expected return, variance and standard deviation for the S&P/TSX Composite Index(illustrate your solution using an embedded Excel Spreadsheet).

c.Using an embedded Excel Spreadsheet, calculate the Covariance and Correlation Coefficient between Lidar’s returns and the market index returns.

d.Calculate the Ex Ante Beta Coefficient for Lidar. Illustrate the formula and your calculations using MS Equation.

e.Is Lidar an aggressive or defensive investment? Explain.

f.Calculate the required rate of return on Lidar stock.

g.Is Lidar over priced, under priced or fairly valued? Explain.

Table 1.1 - Forecast Possible Returns on Lidar Corp. and S&P/TSXPossible

Returns on S&P/TSX

                                                                   Possible Returns          Composite

State                                                  Probability              on Lidar Corp                 Index

Economic Boom                           25%                            47%                           18%

Moderate Growth                        30%                            23%                           11%

Normal                                           25%                            90%                            7%

Recession                                       20%                           -45%                           -6%
Answer #1
Expected Return =Mean Return =SUMof ((Probability)*(Return))
Variance of Return =Sum of(Probability* (Deviation ^2))
Deviation =Return -Mean Return
Standard Deviation of Return =Square Root of Variance of Return
ANALYSIS OF MARKET RETURN(S&P/TSX)
p R1 A1=R1*P D1=R1-8.35 E1=(D1^2) F1=p*E1
State of Economy Probability Return of Lidar(%) Probability*Return(%) Deviation(%) Deviation Squared(%%) Probability*Deviation Squared(%%)
Boom 0.25 18 4.5 9.65 93.1225 23.280625
Moderate Growth 0.3 11 3.3 2.65 7.0225 2.10675
Normal Growth 0.25 7 1.75 -1.35 1.8225 0.455625
Recession 0.2 -6 -1.2 -14.35 205.9225 41.1845
SUM 8.35 SUM 67.0275
Expected Return =Mean return S&P/TSX 8.35 %
Variance of Return 67.0275 %%
Standard Deviation of Return =SQRT(67.0275)= 8.19 %
ANALYSIS OF RETURNS OF LIDAR LIMITED
p R2 A2=R2*p D2=R2-32.15 E2=(D2^2) F2=p*E2
State of Economy Probability Market Return(%) Probability*Return(%) Deviation(%) Deviation Squared(%%) Probability*Deviation Squared(%%)
Boom 0.25 47 11.75 14.85 220.5225 55.130625
Moderate Growth 0.3 23 6.9 -9.15 83.7225 25.11675
Normal Growth 0.25 90 22.5 57.85 3346.6225 836.655625
Recession 0.2 -45 -9 -77.15 5952.1225 1190.4245
SUM 32.15 SUM 2107.3275
Expected Return =Mean return of Lidar Limited 32.15 %
Variance of Return 2107.3275 %%
Standard Deviation of Return =SQRT(2107.3275)= 45.91 %
COVARIANCE BETWEEN RETURNS OF LIDAR AND MARKET(S&P/TSX)
Covariance (A,B)=SUM of (Probability*Deviation of A* Deviation of B)
p D1=R1-8.35 D2=R2-32,15 G=p*D1*D2
State of Economy Probability Deviation of Market(%) Deviation of Lidar (%) probability*Deviation A*Deviation B
Boom 0.25 9.65 14.85 35.825625
Moderate Growth 0.3 2.65 -9.15 -7.27425
Normal Growth 0.25 -1.35 57.85 -19.524375
Recession 0.2 -14.35 -77.15 221.4205
SUM 230.4475
Covariance of return of Lider and Market(S&P/TSX) 230.4475 %%
Correlation of Return A and B=(Covariance)/(Std DeviationA* Standard Deviation B)
Covariance(Lider, market)=230.4475
Correlation of Return Lider and Market=(230.4475/8.19*45.91)= 0.613168
ANSWER:
a Expected Return of Lidar 32.15%
b S&P/TSX Composite Index
Expected Return 8.35%
Variance 67.0275 %%
Standard Deviation 45.91%
c Covariance between return of Lidar and market 230.4475 %%
Correlation between return of Lidar and market 0.61316799
d Beta Coefficient of Lidar
Beta=Covariance/Variance of market=230.4475/67.0275 3.43810376